can you explain pythagorean theorem, step by step thanks
a=8
b=?
c=2rad97

Answer by Theo(13342)

the pythagorean theorem is normally stated as c^2 = a^2 + b^2

a is one leg of a right triangle.
b is the other leg of the right triangle.
c is the hyopenuse of the right triangle.

you are given that a = 8, b = ? c = 2rad97 which i assume is 2 * sqrt(97).

in the equation of c^2 = a^2 + b^2, you can solve for b to get:

b^2 = c^2 - a^2.

you know the value of a and you know the value of c, so you can replace a and c in the formula with those values to get:

b^2 = (2 * sqrt(97))^2 - 8^2

simplify this to get b^2 = 388 - 64

simplify further to get b^2 = 324

take the square root of both sides of this equation to get b = 18

your triangle now becomes:

a=8
b=18
c=2rad97

if this is true, then c^2 must be equal to a^2 + b^2 because pythagorus proved that it had to be way back in the days of ancient greece.

so (2 * sqrt(97))^2 must be equal to 8^2 + 18^2 if we calculated b correctly.

8^2 + 18^2 = 64 + 324 = 388

(2 * sqrt(97))^2 is also equal to 388.

the equation is confirmed to be true and your solution is that the value of b is 18.

here's a reference regarding the pythagorean theorem.

https://www.mathsisfun.com/pythagoras.html